Use a tableau to test this argument for logical correctness.

1. Anyone who understands human nature is clever.

2. Noone is a true poet unless he can stir the hearts of men.

3. Shakespeare wrote Hamlet.

4. Noone who doesn’t understand human nature can stir the hearts of men.

5. Noone but a true poet could have written Hamlet.

6. Therefore, Shakespeare is clever. Use the following propositional language.

a = able to stir the hearts of men c = clever s = Shakespeare p = true poet u = understands human nature w = writer of Hamlet